f(x) = x+3; g(x) = x squared + 1. what is f(g(1)) ?
Answer by stanbon(75887)

f(x) = x+3; g(x) = x squared + 1. what is f(g(1)) ?
-----
f(g(1)) = f(1^2+1) = f(2) = 2+3 = 5
===============
Cheers,
Stan H.